# Recent Advances in Attack and Defense Approaches of Large Language Models

### v2

@@ 4.4 Post-Safety Alignment
A significant branch of post-safety alignment techniques is machine unlearning (Si et al., 2023; Zhang et al., 2024a; Liu et al., 2024a). This paradigm involves selectively forgetting or erasing undesirable knowledge in LLMs, e.g., copyrighted and user privacy content, expecting to mitigate risks associated with model outputs that are potentially influenced by biased or sensitive information. Under this principle, Yao et al. (2023b) have pioneered a work that uses only negative examples to unlearn LLMs, which is done by applying gradient ascent on the loss function of those undesirable samples. Subsequent studies (Lu et al., 2024; Zhang et al., 2024b) have improved this approach by emphasizing the need to retain general knowledge while unlearning harmful knowledge. Instead of tuning all model parameters for unlearning, Chen & Yang (2023) introduce an efficient framework to eliminate the effect of unwanted data by designing separate lightweight unlearning layers. These layers learn to forget different sets of data under the guidance of a selective teacher-student objective. Moreover, Liu et al. (2024f) present a two-stage unlearning framework based on the concept of first isolating and then removing harmful knowledge in model parameters. This framework has been shown to effectively balance the trade-off between removing undesirable information and preserving utility.
Another line of research seeks to prevent LLMs from generating harmful content or resisting jailbreak attacks via steering the decoding process of LLMs. Prompt engineering (White et al., 2023; Chen et al., 2023; Phute et al., 2023; Suo, 2024), as a way to indirectly control the decoding of LLMs, has been preferred due to its ease of operation. To illustrate, Xie et al. (2023) present a simple yet effective technique to defend against various jailbreak attacks by encapsulating the user’s query in a system prompt that reminds ChatGPT to respond responsibly. Further, Wei et al. (2023b) explore a strategy called "In-Context Defense" that teaches the LLM to resist jailbreaking by imitating a few examples of refusing harmful queries. On the other hand, directly manipulating the probability of generated tokens provides a more precise manner to bootstrap the LLM’s decoding process. For instance, Zhong et al. (2024) propose a straightforward method based on the principle of contrastive decoding, which aims to boost the probability of desired safe outputs by suppressing undesired outputs. Besides, Xu et al. (2024a) develop a safety-aware decoding strategy to protect LLMs from jailbreak attacks. This strategy identifies safety disclaimers and amplifies their token probabilities while attenuating the probabilities of token sequences aligned with jailbreak attacks’ objectives.
Pain Point: Although unlearning-based methods succeed in enhancing the safety of the backbone model, they also significantly exacerbate the issue of over-safety. Additionally, utilizing general data for distillation not only incurs extra training costs but also has limited effectiveness in maintaining utility. On the other hand, decoding-based approaches do not address the core problem of harmful output from LLMs, since the harmful knowledge remains within the model. Additionally, they also introduce extra costs during inference. As a result, safety enhancement made by existing approaches typically comes at the cost of significantly increasing over-safety and compromising utility, with the drawbacks outweighing the benefits.
Future Work: As discussed above, future research might have to consider how to effectively achieve the three objectives simultaneously, i.e., safety enhancement, over-safety mitigation, and utility preservation. A recent work by Zhao et al. (2024c) has made a beneficial exploration toward this direction, but it is limited to small-scale LLMs (no more than 13B) and focuses on general safety issues. Therefore, we urge that more efforts should be devoted to developing post-safety alignment techniques that can alleviate over-safety and maintain utility, as well as the comprehensive evaluation of these techniques in a wider range of domains.
@@ 4.5 Model Merge
The methodology of model merging originates from the observation that different fine-tuned models initialized from the same pre-trained backbone model may share a part of the optimization trajectory and diverge on only a fraction of the model parameters oriented to different learning tasks. The parameters of the finely tuned models adapted to different tasks can be hence merged via arithmetic averaging to reach better generalization over out-of-domain input and deliver multi-task learning at the same time. In fact, the idea of merging model parameters to mitigate conflicts between different tasks has been employed and proved to be effective in Federated Learning and continual learning methods. Following the spirit, Zhao et al. (2024b); Kadhe et al. (2024) adopt the model merging methods to achieve a balance between unlearning the unsafe response, yet avoiding over-defensiveness as much as possible. They first assume to have a collection of input questions and well-tagged harmful response texts. They use gradient ascent using the harmful data collection over the backbone model to first compute the model parameter update of the backbone model, dedicated to preventing harmful answers. Furthermore, Zhao et al. (2024b) choose to perform gradient descent over the unaligned model using the harmful input-answer pairs to compute the model parameter update to mitigate over-defense. The two patches of model updates are then integrated using the model merging technique to derive a safety-aligned model balancing the safety alignment and utility.
@@
This line of model merging techniques (Yu et al., 2024; Yadav et al., 2023; Zhang et al., 2023), though operationally simple, still lack deep theoretical investigation regarding two perspectives. First of all, it is unclear how the adversarially updated model parameters with the unlearning objective and tagged harmful responses are associated with the embeddings with safe response. It is possible that adversarial training with a limited set of harmful response texts is prone to overfitting and makes the model still vulnerable to further new jailbreaking prompts. Second, it is difficult to control the over-defensiveness of the merged model parameters. Merging the model parameters does not provide an explicit explanation of how the overdefensive response may be prevented.
@@ 5 Discussion
The field of large language models (LLMs) is indeed dynamic and rapidly evolving, with continuous advancements in both attacks and defenses. As attackers uncover new vulnerabilities and develop sophisticated exploits, defenders are required to respond with equally innovative countermeasures. This ongoing interaction between attackers and defenders is crucial for developing safe and reliable LLM systems.
Measuring attack effectiveness on LLMs is often more straightforward than evaluating the robustness of defense. Nevertheless, the effectiveness of defenses must be scrutinized carefully to avoid a false sense of security. Guaranteeing the defense strength is necessary to ensure that the defenses are not only theoretically sound but also practically effective against real-world threats in LLM systems.
Theoretical analysis of LLMs has indeed struggled to keep pace with their rapid growth and deployment. The sheer complexity and scale of these models pose significant challenges for theoretical frameworks that aim to understand and predict their behavior. While theoretical analysis provides valuable insights, it often falls short in capturing the nuanced performance of LLMs in practical applications.
Recent research has made strides by focusing on scaled case studies and examining specific features and inner mechanisms of LLMs (Lee et al., 2024; Allen-Zhu & Li, 2024; Templeton, 2024). These studies enhance our understanding of model behavior and contribute to improving model interpretability. For instance, the work inspired by (Xu et al., 2024c) emphasizes the need to explore fundamental questions about LLMs, such as the possibility of completely eliminating non-robust features.

### v3

@@ 4.5 Safety Fine-Tuning
Additionally, Rosati et al. (2024) offer a defense mechanism that prevents LLMs from being maliciously fine-tuned for harmful purposes. This mechanism works by removing information about harmful representations such that it is difficult to recover them during fine-tuning. This coincides with the idea of machine unlearning which involves selectively forgetting or erasing undesirable knowledge in LLMs, e.g., copyrighted and user privacy content, expecting to mitigate risks associated with model outputs that are potentially influenced by biased or sensitive information. Under this principle, Yao et al. (2023) have pioneered a work that uses only negative examples to unlearn LLMs, which is done by applying gradient ascent on the loss function of those undesirable samples. Subsequent studies (Lu et al., 2024; Zhang et al., 2024) have improved this approach by emphasizing the need to retain general knowledge while unlearning harmful knowledge. Instead of tuning all model parameters for unlearning, Chen & Yang (2023) introduce an efficient framework to eliminate the effect of unwanted data by designing separate lightweight unlearning layers. These layers learn to forget different sets of data under the guidance of a selective teacher-student objective. Moreover, Liu et al. (2024e) present a two-stage unlearning framework based on the concept of first isolating and then removing harmful knowledge in model parameters. This framework has been shown to effectively balance the trade-off between removing undesirable information and preserving utility.
Apart from the aforementioned defenses, recent works also focus on improving the safeguards of open-weight models, where adversaries have full access to the model weights and can tamper with built-in safeguards. To address this issue, Tamirisa et al. (2024) propose a novel method called Tampering Attack Resistance (TAR), which consists of two phases: model safeguarding and tamper-resistance training. During the model safeguarding phase, an initial safeguard is applied to the base model to restrict access to harmful knowledge or behaviors. In the tamper-resistance training phase, the model is trained against a set of tampering attacks using a novel adversarial training procedure. This procedure aims to maximize a proxy safety metric after applying an adversarial attack to the model. By doing so, the TAR method ensures that the safeguards cannot be easily removed by adversaries, even after thousands of steps of fine-tuning.
Challenges: There is a well-known trade-off in enhancing LLM’s instruction-following capabilities while ensuring they remain safe and reliable. While integrating safety-related examples into the fine-tuning process can bolster a model’s safety features, this could inadvertently stifle its adaptability to benign inputs or even compromise its performance across certain tasks. Moreover, the reliance on manual curation of safety-related examples or prompt templates demands considerable human expertise and is prone to subjective bias, which can affect consistency and reproducibility across different applications. On the other hand, fine-tuning itself can be used both to enhance the safety of LLMs by adding safety-related examples, and to attack LLMs by introducing adversarial examples. The mechanism proposed by Rosati et al. (2024), while seemingly promising, is limited to defending against supervised fine-tuning attacks in LLMs. Also, it requires paired safe and unsafe examples, which makes data collection more expensive and complex.
Future Work: Future work on safety fine-tuning perhaps should strive to achieve a win-win situation for both the safety and utility of LLMs. One promising avenue lies in the advancement of automated systems for identifying and incorporating safety signals without presupposing extensive domain knowledge. Integrating reinforcement learning strategies could potentially help models dynamically adjust their responses based on feedback, thereby improving both safety and task-specific performance. At the same time, future research should invest in stronger attack settings to emulate worst-case attacks during the fine-tuning process and investigate different types of harm, based on which more effective and comprehensive defense mechanisms need to be developed.
@@ 4.6 Model Merge
Model merge is initially used to combine the capabilities of multiple models, enhance model robustness, and improve out-of-distribution performance (Cao et al., 2022; Wortsman et al., 2022). Recently, model merge also represents a valuable technique within the arsenal of defenses (Yang et al., 2024). It is not a standalone solution but rather a complementary method that, when used with other defenses, can significantly enhance the robustness of LLMs. By leveraging the diversity of multiple models, merging can provide a more resilient system that is better equipped to withstand adversarial manipulations.
@@
The methodology of model merging originates from the observation that different fine-tuned models initialized from the same pre-trained backbone model may share a part of the optimization trajectory and diverge on only a fraction of the model parameters oriented to different learning tasks. The parameters of the finely tuned models adapted to different tasks can be hence merged via arithmetic averaging to reach better generalization over out-of-domain input and deliver multi-task learning at the same time. In fact, the idea of merging model parameters to mitigate conflicts between different tasks has been employed and proved to be effective in federated learning and continual learning methods. Following the spirit, Zhao et al. (2024a) and Kadhe et al. (2024) adopt the model merging methods to achieve a balance between unlearning the unsafe responses, yet avoiding over-defensiveness as much as possible. They first assume that they have a collection of input questions and well-tagged harmful response texts. They use gradient ascent using the harmful data collection over the backbone model to first compute the model parameter update of the backbone model, dedicated to preventing harmful answers. Furthermore, Zhao et al. (2024a) choose to perform gradient descent over the unaligned model using the harmful input-answer pairs to compute the model parameter update to mitigate over-defense. The two patches of model updates are then integrated using the model merging technique to derive a safety-aligned model balancing the safety alignment and utility.  Liu et al. (2024a) demonstrate the assistance of contaminated pre-trained models through the merging of a defensive LoRA model (Hu et al., 2021). The experimental analysis by  Gallego (2024) indicates that model merging offers an effective defense mechanism against jailbreak attacks.
@@
Challenges: This line of model merging techniques (Yu et al., 2024; Yadav et al., 2024; Zhang et al., 2023b), though operationally simple, still lack deep theoretical investigation regarding two perspectives. First of all, it is unclear how the adversarially updated model parameters with the unlearning objective and tagged harmful responses are associated with the embeddings with safe responses. It is possible that adversarial training with a limited set of harmful response texts is prone to overfitting and makes the model still vulnerable to further new jailbreaking prompts. Second, it is difficult to control the over-defensiveness of the merged model parameters. Merging the model parameters does not provide an explicit explanation of how the overdefensive response may be prevented.
Future Work: To address the current theoretical gaps and practical limitations, future research can delve into the intricate relationship between adversarially updated model parameters, unlearning objectives, and the embeddings of safe responses. Understanding this connection is crucial for developing robust models that can withstand novel jailbreaking prompts without overfitting to a limited set of harmful response texts. Additionally, controlling the over-defensiveness of merged model parameters is paramount. Future work should focus on creating explicit mechanisms to prevent overdefensive responses, ensuring that the merged models strike a balance between security and usability. This may involve developing new merging strategies that incorporate explainability and interpretability, allowing for better oversight and fine-tuning of the model’s defensive behavior. By tackling these challenges, the model merge technique can evolve into a more reliable and effective defense against a wide range of adversarial attacks.
5 Related Work
LLMs have significantly advanced natural language processing, but their integration into various domains has introduced substantial security and privacy challenges. Recent surveys have extensively explored these aspects, shedding light on both the benefits and risks associated with LLMs. There are also surveys that focus more on investigating specific attack and defense techniques.
LLM security and privacy challenges.
Yao et al. (2024) investigate LLMs’ impact on security and privacy, categorizing research into beneficial applications, offensive uses, and inherent vulnerabilities, emphasizing their potential to enhance code security and data privacy while also being susceptible to user-level attacks. Chowdhury et al. (2024) analyze attacks targeting LLMs, including adversarial attacks, data poisoning, and privacy concerns, discussing their mechanisms, impacts, and current defense strategies to foster awareness and inspire robust solutions. Das et al. (2024) review LLM security and privacy challenges, including jailbreaking, data poisoning, and Personally Identifiable Information (PII) leakage attacks, assessing vulnerabilities, investigating emerging attacks, and reviewing potential defense mechanisms, outlining research gaps and future directions. Dong et al. (2024) offer a comprehensive overview of recent studies on LLM conversation safety, covering attacks, defenses, and evaluations. Chua et al. (2024) survey recent AI safety research trends for Generative AI LLMs, emphasizing the need for unified theories to address safety challenges and encouraging the development of aligned and secure models.
LLM attacks and defenses.
Two recent studies extensively analyze jailbreaking in LLMs. The survey of Xu et al. (2024b) evaluate nine attack techniques and seven defense techniques across three language models, revealing the underperformance of existing white-box attacks and the significant impact of special tokens on attack success. While the work of Yi et al. (2024a) proposes a detailed taxonomy of jailbreak attack and defense methods, categorizing attacks into black-/white-box based on model transparency and defenses into prompt-/model-level, with investigating current evaluation methods. Liu et al. (2024d) formalize prompt injection attacks, design a new attack by combining existing ones, and establish a common benchmark for future research by conducting a systematic evaluation of five attacks and ten defenses across ten LLMs and seven tasks. Huang et al. (2024c) address the safety risks of harmful fine-tuning attacks in the fine-tuning-as-a-service model, aiming to clarify misunderstandings and establish the research problem formally, while outlining future research directions.
Our survey aims to provide a structured summary that enhances understanding of LLM conversation safety and encourages further investigation into this critical subject. We provide a holistic review of LLM vulnerabilities, threats, and defense mechanisms, analyzing recent studies on attack vectors and model weaknesses. We offer insights into attack mechanisms and the evolving threat landscape. By contrasting advancements in attack and defense methodologies, our work identifies research gaps and proposes future directions to enhance LLM security, advancing the understanding of LLM safety challenges and guiding the development of more robust security measures.
@@ 6 Discussion
The field of LLMs is indeed dynamic and rapidly evolving, with continual advancements in both attack methodologies and defense mechanisms. As attackers uncover new vulnerabilities and develop sophisticated exploits, defenders are compelled to respond with equally innovative countermeasures. This ongoing interaction between attackers and defenders is crucial for developing safe and reliable LLM systems.
Measuring attack effectiveness on LLMs is often more straightforward than evaluating the robustness of defenses. While attack success can be quantified using specific metrics such as the rate of undesired outputs or successful policy evasion, assessing the strength of defenses requires careful scrutiny to avoid a false sense of security. Moreover, evaluation metrics for both attacks and defenses should take practical usage scenarios into consideration. For instance, safety alignment should consider continual fine-tuning on specialized datasets by end-users and how it affects previous alignments. By recognizing that both attacks and defenses operate within a dynamic environment, researchers can incorporate longitudinal studies and simulations that mimic real-world usage and adaptation.
Theoretical analysis of LLMs has indeed struggled to keep pace with their rapid growth and deployment. The sheer complexity and scale of these models pose significant challenges for theoretical frameworks that aim to understand and predict their behavior. While theoretical analysis provides valuable insights, it often falls short of capturing the nuanced performance of LLMs in practical applications.
Recent research has made strides by focusing on scaled case studies and examining specific features and inner mechanisms of LLMs (Lee et al., 2024; Allen-Zhu & Li, 2023; Templeton et al., 2024). These studies enhance our understanding of model behavior and contribute to improving model interpretability. For instance, the findings of Wei et al. (2024a) underscore the necessity of safety-capability parity, where safety mechanisms should be as sophisticated as the underlying model. This parity is essential to prevent attacks from exploiting cutting-edge capabilities of the model that less advanced safety mechanisms cannot detect or address. The work of Xu et al. (2024c) emphasizes the need to explore fundamental questions about LLMs, such as the possibility of completely eliminating non-robust features.